### v1

# State-action control barrier functions: Imposing safety on learning-based control with low online computational costs

Abstract
Learning-based control with safety guarantees usually requires real-time safety certification and modifications of possibly unsafe learning-based policies. The control barrier function (CBF) method uses a safety filter containing a constrained optimization problem to produce safe policies. However, finding a valid CBF for a general nonlinear system requires a complex function parameterization, which in general, makes the policy optimization problem difficult to solve in real time. For nonlinear systems with nonlinear state constraints, this paper proposes the novel concept of state-action CBFs, which not only characterize the safety at each state but also evaluate the control inputs taken at each state. State-action CBFs, in contrast to CBFs, enable a flexible parameterization, resulting in a safety filter that involves a convex quadratic optimization problem. This, in turn, significantly alleviates the online computational burden. To synthesize state-action CBFs, we propose a learning-based approach exploiting Hamilton-Jacobi reachability. The effect of learning errors on the effectiveness of state-action CBFs is addressed by constraint tightening and introducing a new concept called contractive CBFs. These contributions ensure formal safety guarantees for learned CBFs and control policies, enhancing the applicability of learning-based control in real-time scenarios. Simulation results on an inverted pendulum with elastic walls validate the proposed CBFs in terms of constraint satisfaction and CPU time.
Index Terms:
Constrained control, control barrier functions, machine learning, nonlinear control.
I Introduction
Learning-based control methods have demonstrated extensive success across many applications, such as autonomous vehicles [1] and robot manipulation [2]. Both supervised learning and reinforcement learning (RL) have been widely used for controller synthesis. However, learning-based controllers may provide unsafe control actions that result in undesirable or even destructive effects on the system. In control systems, safety means that the trajectories of the closed-loop system should satisfy state and input constraints for the entirety of the system’s evolution. The lack of safety guarantees limits the ability of learning-based control to achieve safety-critical tasks.
There has been an increasing interest in designing strategies to make learning-based controllers safe. Among the available methodologies including penalty methods [3], constraint elimination [4], and using invariant sets [5, 6], using control-barrier functions (CBFs) [7, 8] for safe controller synthesis is a universal and convenient approach. CBFs are energy-based functions and use their sub-level sets to characterize the safety of the dynamical system, avoiding the use of complex invariant set representations. Given any learning-based controller without safety considerations, an online optimization problem, called a safety filter, is then solved to find the closest safe input to this controller, while safety is gained by adding a constraint based on a CBF into the optimization problem.
In the realm of CBF-based methodologies, a prominent limitation is the absence of universally applicable techniques for generating valid CBFs, thereby necessitating reliance on manually designed or problem-specific CBFs. For some particular kinds of systems such as linear, piecewise affine [9], or polynomial systems [10], CBFs can be computed by solving convex optimization problems. For nonlinear systems with general constraints, using learning-based algorithms accompanied by advanced function approximators to estimate CBFs has been explored in several contexts. Roughly, there are four kinds of methods to learn CBFs: the optimization-based method [11], the learner-verifier method [12], the Hamilton-Jacobi (HJ) reachability method [13], and the predictive safety filter method [14, 15]. The first two methods can be conservative, i.e., the learned safe set can be a small subset of the maximal controlled invariant set. In comparison, the HJ reachability method relies on the dynamic programming principle to approximate the maximal controlled invariant set iteratively, while the predictive safety filter method is based on a receding-horizon open-loop optimal control problem to implicitly determine the safe set, which also converges to the maximal controlled invariant set as the horizon goes to infinity [16]. A comprehensive comparison of HJ reachability, predictive safety filters, and CBFs can be found in [17].
No matter how the CBF is learned, to control a given system with safety guarantees, it is inevitable to solve an online optimization problem with CBF-based constraints, which is usually non-convex for discrete-time nonlinear systems [18]. To obtain a satisfactory approximation accuracy, sophisticated function approximators such as deep neural networks (NNs) are commonly used to parameterize the CBF [14, 7]. However, this will inherently cause non-convexity and increase the complexity of the optimization problem. As a result, the increased online computational load makes the CBF-based approach unsuitable for situations where fast computation of control inputs is required. Besides, the effect of approximation errors on the validity of the learned CBFs and the resulting control policies has not yet been addressed, as highlighted in [7].
This paper contributes to the state of the art in the following aspects. To impose safety on nonlinear systems at low online computational costs, we propose the novel concept of state-action CBFs, the zero sub-level sets of which can be used in the safety filter to generate a safe policy. Compared to standard CBFs, state-action CBFs allow for a flexible parameterization so that the resulting safety filter contains a convex quadratic optimization problem. Consequently, the online computation burden is greatly eased. To analyze the effect of approximation errors on learned CBFs, we propose a constraint tightening approach and introduce a novel concept of contractive CBFs. When learning such CBFs, the invariance property of CBFs can be retained, despite the presence of approximation errors. We also explore the connection between state-action CBFs and (contractive) CBFs, and develop a new learning-based method to approximate state-action CBFs. As a result, the safety of the policy filtered by state action CBFs is preserved for a sufficiently small approximation error.
II Preliminaries and problem formulation
II-A Preliminaries
We consider a deterministic discrete-time nonlinear system
$x_{t+1}=f\left(x_{t},u_{t}\right),\quad t=0,1,\ldots,$ (1)
where $x_{t}\in\mathcal{X}\subseteq\mathbb{R}^{n_{x}}$ and $u_{t}\in\mathcal{U}\subseteq\mathbb{R}^{n_{u}}$ are the state and the input at time step $t$, and $f(\cdot,\cdot):\mathcal{X}\times\mathcal{U}\to\mathcal{X}$ is a continuous function satisfying $f(0,0)=0$. We consider a constrained optimal control problem in which the states and inputs should satisfy time-invariant constraints: $x_{t}\in X:=\{x\in\mathcal{X}|h(x)\leq 0\}$ and $u_{t}\in U\subseteq\mathcal{U}$. Here, $h(\cdot):\mathcal{X}\to\mathbb{R}$ is a continuous function that defines the state constraint [^1] For the constraint defined by multiple inequalities $h_{i}(x)\leq 0,\;i=1,2,...,I$, we can let $h(x)=\max_{i\in\{1,...,I\}}h_{i}(x)$. The set $\{x|h(x)\leq 0\}$ is then identical to $\{x|h_{i}(x)\leq 0,\;i=1,2,...,I\}$, and $h$ will be continuous if each $h_{i}$ is continuous.. Throughout the paper, we assume that we completely know $f$. Besides, we assume that $X$ is compact and that $U$ is a polytope.
The control objective is to regulate a redefined control policy, e.g., a policy learned by reinforcement learning or a policy that approximates model predictive control (MPC) [19], so that the safety of the system is ensured.
II-B Control barrier functions
To achieve the objective, we need to design a control policy that can ensure constraint satisfaction at all time steps. A set of states for which such a policy exists, needs to be defined. Usually, this class of sets is called controlled-invariant sets or safe sets. For high-dimensional systems, however, some controlled-invariant sets could have complex representations, making the controller synthesis difficult. A CBF uses the sub-level set of a scalar function to conveniently define the safe set.
Definition 1 (Control barrier function)
A continuous function $B(\cdot):\mathcal{X}\to\mathbb{R}$ is called a control barrier function (CBF) with a corresponding safe set $\mathcal{S}_{B}:=\left\{x\in\mathbb{R}^{n_{x}}|B(x)\leq 0\right\}\subseteq\mathcal{X}$, if $\mathcal{S}_{B}$ is non-empty,
$h(x)\leq 0,\;\forall x\in\mathcal{S}_{B},$ (2)
and
$\min_{u\in U}B(f(x,u))\leq 0,\;\forall x\in\mathcal{S}_{B}.$ (3)
Furthermore, a CBF is called an exponential CBF if it satisfies (2) with 0 replaced by $B(x)$ and if there exists a $\beta\in(0,1)$ such that
$\min_{u\in U}B(f(x,u))\leq\beta B(x),\;\forall x\in\mathcal{S}_{B}.$ (4)
Condition (2) is equivalent to $\mathcal{S}_{B}\subseteq X$, which is usually assumed in literature [18]. With a CBF available, one generate a safe control policy in $\mathcal{S}_{B}$ by using the following optimization-based approach:
$\pi_{\mathrm{safe}}(x)=\arg\min_{u\in U}$ $\;||u-\pi_{0}(x)||_{2}$
[…]
which serves as a safety filter [5] for any unsafe policy $\pi_{0}$. The definition of a safe policy is formally given as follows.
Definition 2
A policy $\pi(\cdot):\mathcal{X}\to\mathcal{U}$ is safe in $\mathcal{S}\subseteq X$ for system (1) under the state and input constraints $x\in X,\;u\in U$, if $\pi(x)\in U,\;\forall x\in\mathcal{S}$, and for any initial state in $\mathcal{S}$, the state trajectory of (1) steered by $\pi$ will always stay in $\mathcal{S}$.
For general nonlinear systems with state and input constraints, synthesizing a non-conservative CBF is a difficult task. To deal with this issue, using advanced function approximators to learn a CBF certificate has received much attention. One can see [7] for a comprehensive survey.
II-C Problem formulation
Problem P1: high online computational complexity. One main limitation of (II-B) is that it needs to solve a usually non-convex optimization problem in real time. If a CBF is represented by a complex function approximator $B_{\theta}(\cdot)$ with the parameter $\theta$, solving (II-B) may take much online computation time and result in very sub-optimal solutions. Even if a convex $B_{\theta}$ is formed, the constraint in (II-B) could be non-convex due to the nonlinearity of $f$.
Problem P2: effects of approximation errors. Another limitation of (II-B) is that the approximation error of $B(\cdot)$ may affect the safety of the system controlled by the optimizer of (II-B) with $B(\cdot)$ replaced by $B_{\theta}(\cdot)$. Besides, the recursive feasibility of (II-B) is also not guaranteed with $B_{\theta}(\cdot)$.
III State-action control barrier function
To deal with Problem P1, motivated by Q-learning in RL [20], we propose a novel safety filter in the following form:
$\pi_{\mathrm{safe}}(x)=\arg\min\{||u-\pi_{0}(x)||_{2},\;\mathrm{s.t.}\;u\in U,\;Q(x,u)\leq 0\},$ (6)
where $Q(\cdot,\cdot):\mathcal{X}\times U\to\mathbb{R}$ is a function of states and actions. We will analyze how to enforce the safety of $\pi_{\mathrm{safe}}$ by imposing conditions on $Q$. To achieve this, we introduce the definition of state-action CBFs as follows.
Definition 3 (State-action control barrier function)
A continuous function $Q(\cdot,\cdot):\mathcal{X}\times U\to\mathbb{R}$ is called a state-action control barrier function with a corresponding safe set $\mathcal{S}_{Q}:=\{x\in X|\min_{u\in U}Q(x,u)\leq 0\}$, if $\mathcal{S}_{Q}$ is non-empty, and for any $x\in\mathcal{S}_{Q}$, any $u\in U$ that satisfies $Q(x,u)\leq 0$ will make $f(x,u)\in\mathcal{S}_{Q}$.
With the definition of state-action CBF, we have the following properties:
Lemma 1
(i) Consider the safety filter (6), any state-action CBF $Q$ will render $\pi_{\mathrm{safe}}$ safe in $\mathcal{S}_{Q}$. In other words, $\mathcal{S}_{Q}$ is control-invariant. (ii) If $B$ is a CBF, $Q(\cdot,\cdot)=B(f(\cdot,\cdot))$ will be a state-action CBF and the safe sets satisfy $\mathcal{S}_{B}\subseteq\mathcal{S}_{Q}$.
The proof of these properties is given in Appendix -A.
Similarly to designing CBF, the challenge of using (6) is that the explicit form of a state-action CBF cannot be directly obtained based on Definition 3. However, as we know that $Q(\cdot,\cdot)=B(f(\cdot,\cdot))$ is a state-action CBF if $B$ is a CBF, if the exact value $B(x^{\mathrm{s}})$ of a CBF on any state sample $x^{\mathrm{s}}$ is available, we can use supervised learning-based methods to approximate $Q$, using a parameterized function approximator $Q_{\theta}$ with the parameter $\theta$.
The advantage of directly approximating $Q$ over approximating $B$ is that we can design a specific structure for $Q_{\theta}$ to simplify the constraint in (6), so that the online computational cost of solving (6) is reduced. To achieve this, we can specify the following parameterization:
$Q_{\theta}(x,u)=q_{1,\theta}(x)+q_{2,\theta}(x)u+u^{T}Q_{3,\theta}(x)u,\;\mathrm{with}\;Q_{3,\theta}\succeq 0,$ (7)
which will make problem (6) a convex QP with linear and convex quadratic constraints. In (7), $q_{1,\theta}(\cdot):\mathbb{R}^{n_{x}}\to\mathbb{R}$, $q_{2,\theta}(\cdot):\mathbb{R}^{n_{x}}\to\mathbb{R}^{n_{x}}$, and $Q_{3,\theta}(\cdot):\mathbb{R}^{n_{x}}\to\mathbb{R}^{n_{x}\times n_{x}}$ are parameterized functions with all parameters condensed in $\theta$.
IV Constraint tightening and contractive CBF
To cope with Problem P2, we propose a method to compute conservative CBFs based on state constraint tightening. In particular, we consider two kinds of conservative CBFs. The first kind includes the CBFs of system (1) under the tightened state constraint $X_{\lambda}:=\{x\in\mathcal{X}|h(x)+\lambda\leq 0\}$ where $\lambda$ is a positive constant. In this situation, the CBFs still satisfy (3) or (4), while condition (2) is tightened to $h(x)+\lambda\leq B(x),\forall x\in\mathcal{S}_{B}$.
The second one, called the contractive CBF, is defined as follows
Definition 4 (Contractive CBF)
A CBF $B(\cdot):\mathcal{X}\to\mathbb{R}$ is called a $\lambda$-contractive CBF [^2] Note that in fact the definition of contractive CBFs given here is not consistent with the common definition of contraction mappings [20]. Formally speaking, the CBF we define should be called a CBF with a contractive safe set [21]. However, for compactness we adopt “contractive CBF” in the paper. if there exists a $\lambda>0$ such that
$\min_{u\in U}B(f(x,u))\leq-\lambda,\;\forall x\in\mathcal{S}_{B}.$ (8)
Similarly, an exponential CBF $B(\cdot):\mathcal{X}\to\mathbb{R}$ is called a $\lambda$-contractive exponential CBF if there exist $\lambda>0$ and $\beta\in(0,1)$ such that
$\min_{u\in U}B(f(x,u))\leq\beta B(x)-\lambda,\;\forall x\in\mathcal{S}_{B}.$ (9)
The definition of contractive CBFs is introduced to endow the safe set $\mathcal{S}_{B}$ with a contractive property. It enforces an energy decay when the state is on the boundary of the safe set. In particular, if $B(\cdot)$ is a $\lambda$-contractive CBF, for any $x\in\partial\mathcal{S}_{B}$, there exists an input $u\in U$ such that $B(f(x,u))\leq-\lambda$, which means that $x^{+}=f(x,u)$ is in the interior of $\mathcal{S}_{B}$. Such a property can guarantee that any approximation of $B(\cdot)$ is still a valid CBF if the approximation error is sufficiently small.
V Approximating state-action CBF
In this section, we propose a method for approximating the state-action CBF. Inspired by HJ reachability [13] and predictive CBF [14], we propose a comprehensive framework that uses the optimal value functions of a sequence of optimization problems to implicitly represent the CBF sequence $\{B_{k}\}^{\infty}_{k=1}$. By tuning some parameters, this framework can compute standard CBFs, exponential CBFs, as well as contractive CBFs. Although the exact formulation of each $B_{k}$ is unknown, we can use sampling methods to collect data tuples $(x^{\mathrm{s}},\;u^{\mathrm{s}},\;B_{k}(f(x^{\mathrm{s}},u^{\mathrm{s}}))$, and use these tuples to learn a state-action CBF.
V-A CBF generator based on HJ reachability
We consider each $B_{k}$ as the value function of the following optimization problem:
$B_{k}(x):=$ $\min_{\{x_{t},u_{t}\}_{t=0}^{k}}\max\left\{\max_{t\in\mathbb{I}(k-1)}\alpha^{t}(h(x_{t})+\lambda_{t}),\alpha^{k}B_{0}(x_{k})\right\}$
$\quad\mathrm{s.t.}\;x_{t+1}=f\left(x_{t},u_{t}\right),\;\;u_{t}\in U,\;t=0,1,...,k-1,$
$\quad\quad\quad x_{0}=x.$ (10)
In (V-A), $\mathbb{I}(k-1)$ denotes the set $\{0,1,...,k-1\}$, $\alpha\geq 1$, and $B_{0}$ is a CBF, which, however, could have a very small safe set. We take $\alpha=1$ if $B_{0}$ satisfies (3), or let $\alpha=1/\beta$ if $B_{0}$ is an exponential CBF. To solve (V-A), we need to know the explicit formulation of $B_{0}$. An optimization-based method that can compute a local quadratic $B_{0}$ for the nonlinear system (1) is reported in [15], by linearizing the model around the origin and solving an LMI problem. In Appendix -D, this method is extended to our situations where additional conditions such as (8), (9), and (11)-(12) in the subsequent Theorem 1 are required. Besides, $\lambda_{t}\geq 0,\;t=0,1,...,k-1$ are tuning parameters. We consider the following three options for choosing $\lambda_{t}$:
Option 1: $\lambda_{t}=0$. Option 2: $\lambda_{t}=\lambda$, where $\lambda>0$ is a constant. Option 3: $\lambda_{t}=t\lambda$, where $\lambda>0$ is a constant.
Option 1 means that we are constructing CBFs for the system (1) under the original state constraints. If Option 2 is chosen, it is seen from (V-A) that we increase the function $h$ to $h+\lambda$, i.e., we tighten the original state constraints to $x\in X_{\lambda}=\{x\in\mathcal{X}|h(x)+\lambda\leq 0\}$. More conservatively, in Option 3, we require a linear decrease rate for $h$ w.r.t. the time step to construct contractive CBFs.
In the original HJ reachability analysis [13], $B_{0}(\cdot)$ is chosen as $h(\cdot)$, which is not a CBF. It has been proven in [13] that only when $k=\infty$, $B_{\infty}$ is a CBF. Although this result shows a strong connection between CBFs and HJ reachability, it cannot be used in practice since we cannot solve (V-A) with $k=\infty$. In our case, we require the initial function $B_{0}$ to be a CBF. As a result, any $B_{k},\;k=1,2,...$ will become a CBF, with a non-shrinking safe set as $k$ increases. The following theorem formally states this property, and the proof is given in Appendix -B.
Theorem 1
Consider $B_{k}$ from (V-A). Suppose that $B_{0}$ is a (exponential) CBF for system (1) with the state constraint $x\in X$. We have the following results:
  1. 1.
    Let $\lambda_{t}=0,\;\forall t\in\{0,1,...,k-1\}$. Then $B_{k},\;k=1,2,...$ is a (exponential) CBF for system (1) with the state constraint $x\in X$.
  2. 2.
    Let $\lambda_{t}=\lambda>0,\;\forall t\in\{0,1,...,k-1\}$. If $\lambda,B_{0}$ satisfy
    $h(x)+\lambda\leq B_{0}(x),\forall x\in\mathcal{S}_{B_{0}},$ (11)
    then $B_{k},\;k=1,2,...$ is a (exponential) CBF for system (1) with the tightened state constraint $x\in X_{\lambda}=\{x|h(x)+\lambda\leq 0\}$.
  3. 3.
    Let $\lambda_{t}=t\lambda,\;\forall t\in\{0,1,...,k-1\}$, $\lambda>0$. If $B_{0}$ is a $\lambda$-contractive (exponential) CBF and $k,\lambda,B_{0}$ satisfy
    $h(x)+k\lambda\leq B_{0}(x),\forall x\in\mathcal{S}_{B_{0}},$ (12)
    then $B_{k},\;k=1,2,...$ is a $\lambda$-contractive (exponential) CBF for system (1) with the state constraint $x\in X$.
  4. 4.
    In the statements 1-3, $B_{k}(\cdot)$ is a continuous function in $\mathcal{X}$, and the safe sets satisfy $\mathcal{S}_{B_{0}}\subseteq\mathcal{S}_{B_{1}}\subseteq...\subseteq\mathcal{S}_{B_{\infty}}$. Furthermore, if $f$, $h$, and $B_{0}$ are Lipschitz continuous, $B_{k}$ is Lipschitz continuous.
If $B_{k}$ and $f$ are (Lipschitz) continuous in their domains, then the function $Q$ defined by
$Q(x,u)=B_{k}(f(x,u))$ (13)
is also (Lipschitz) continuous in $\mathcal{X}\times U$.
The condition (12) indicates that the horizon $k$ should not be selected too large when using (V-A) to generate a contractive CBF.
The structure of (V-A) is similar to that in [14, 16]. The main difference is that we use the maximum of $h(\cdot)+\lambda_{t}$ and $B_{0}(\cdot)$ over state trajectories in a finite horizon, while [14, 16] uses the summation of $\max\{h(\cdot),0\}$ and $\max\{B_{0}(\cdot),0\}$ over state trajectories in a finite horizon. This difference means that the safe set $\mathcal{S}_{B_{k}}$ is the zero-sublevel set of $B_{k}$ in our situation, while the safe set in [14, 16] is the zero-level set of $B_{k}$, i.e., the set $\{x\in\mathbb{R}^{n_{x}}|B_{k}(x)=0\}$. Besides, in [14, 16] the weight on the terminal CBF $B_{0}$ needs to be carefully selected.
V-B Learning state-action CBF from CBF samples
After the CBF generator is provided with a fixed $k$, a regression model can be trained to obtain an approximation of a state-action CBF.
First, state and input samples are collected in a compact region $\Omega\times U$ of interest. The region $\Omega$ is task-specified and should be contained in the space where the system is physically realistic. If the state space $\mathcal{X}$ of the system is compact and small, letting $\Omega=\mathcal{X}$ is an ideal choice. If $\mathcal{X}$ is unbounded or too large, one possible choice for $\Omega$ is $\Omega=\{x\in\mathcal{X}|B_{k}(x)\leq\bar{B}\}$ [14], where $\bar{B}>0$ is a positive constant. In the case of $\alpha>1$, the values of $B_{k}$ and $Q$ are likely to grow exponentially as $k$ increases, so it is necessary to specify $\Omega$ as $\{x\in\mathcal{X}|B_{k}(x)\leq\bar{B}\}$ to avoid approximating probably unbounded $Q$. Besides, various sampling methods [22], such as (quasi) random sampling and sampling from a uniform grid, can be applied.
After $N$ samples $\{x^{\mathrm{s}}_{i},u^{\mathrm{s}}_{i}\}^{N}_{i=1}$ are collected, problem (V-A) is solved with $x$ specified as each $f(x^{\mathrm{s}}_{i},u^{\mathrm{s}}_{i})$. As a result, $N$ data tuples $\{(x^{\mathrm{s}}_{i},\;u^{\mathrm{s}}_{i},\;Q(x^{\mathrm{s}}_{i},u^{\mathrm{s}}_{i}))\}^{N}_{i=1}$ are obtained after substituting $x=x^{\mathrm{s}}_{i},\;u=u^{\mathrm{s}}_{i}$ into (13). For general nonlinear systems, (V-A) is usually a nonlinear non-convex optimization problem, which requires a multi-start strategy [23] to find a sufficiently good optimum. For linear systems with linear or ellipsoidal constraints, (V-A) is a convex quadratically constrained quadratic program (QCQP) and the global optimum can be conveniently obtained via gradient-based optimization methods. For piecewise affine systems with piecewise affine constraints, (V-A) can be regarded as a mixed-integer QCQP, which can be globally solved by the branch-and-bound approach [24].
In this work, we use neural networks (NNs) to approximate the state-action CBF. For the parameterization (7), we use three NNs to represent $q_{1,\theta},\;q_{2,\theta}$ and $Q_{3,\theta}$ in (7), respectively. There are several ways to guarantee the positive-semidefiniteness of $Q_{3,\theta}$. For example, we can further parameterize $Q_{3,\theta}$ by $Q_{3,\theta}=L_{\theta}L^{T}_{\theta}$, where $L_{\theta}\in\mathbb{R}^{n_{u}\times n_{u}}$ is a lower triangular matrix with non-negative diagonal entries. Alternatively, we can parameterize $Q_{3,\theta}$ by $Q_{3,\theta}=P_{0}\mathrm{diag}(r_{\theta})\mathrm{diag}(r_{\theta})P^{T}_{0}$, where $r_{\theta}\in\mathbb{R}^{n_{u}}$ is the output of an NN and $P_{0}\in\mathbb{R}^{n_{u}\times n_{u}}$ is a given matrix.
With the specified regression model and the training data, the parameter $\theta$ is optimized to minimize the mean square error of $Q$ over all state-action samples.
V-C Performance analysis
After the approximation $Q_{\theta}$ is obtained, it can be integrated into the safety filter (6), i.e., $Q$ in (6) is replaced by $Q_{\theta}$. In general, the safety of the policy generated from (6) will not be ensured due to the approximation error. However, in this section, we will demonstrate that we can guarantee safety in the presence of a small approximation error. To achieve this, we need an assumption on the boundedness of this error.
Assumption 1
The approximation error is uniformly bounded in $\Omega\times U$. In other words, there exists a non-negative constant $\Delta<\infty$ such that
$|Q_{\theta}(x,u)-Q(x,u)|\leq\Delta,\;\forall(x,u)\in\Omega\times U.$ (14)
Assumption 1 is common in the literature studying performance guarantees of learning-based control [25, 26, 14, 27]. If the chosen function approximator represents a continuous function, since $Q$ is also continuous, the approximation error is always upper bounded in any compact region. To obtain the upper bound, we can first get the error bound on a finite number of samples, and then extract a statistical estimate [25] or compute a deterministic bound in the whole region by using the Lipschitz property of $Q$ and $Q_{\theta}$ [27].
With Assumption 1, we modify the safety filter (6) to
$\pi_{\theta}(x)=\arg\min\{||u-\pi_{0}(x)||_{2},\;\mathrm{s.t.}\;u\in U,\;Q_{\theta}(x,u)\leq\Delta_{i}\},$ (15)
where $i=1,2$, or $3$, corresponding to Options 1,2, and 3 chosen in the CBF generator (V-A), $\Delta_{1}=\Delta_{2}=0$ and $\Delta_{3}=-\lambda+\Delta$.
The following theorem, the proof of which is given in Appendix -C, characterizes the safety performance of the policy $\pi_{\theta}$.
Theorem 2
Consider the system (1) controlled by the policy $\pi_{\theta}$, the state-action CBF $Q$ from (13), and the safety filter (15). Suppose that Assumption 1 is fulfilled. We have the following results: (i) If $B_{k}$ is a CBF for the original state constraint, for any initial state $x_{0}\in\mathcal{S}_{B_{k}}$ that makes problem (15) recursively feasible, the closed-loop system $x_{t+1}=f\left(x_{t},\pi_{\theta}(x_{t})\right),\;t=0,1,\ldots$ has the maximum constraint violation $\Delta$, i.e., $\max_{t\in\mathbb{I}(\infty)}\{h(x_{t})\}\leq\Delta$. (ii) If $B_{k}$ is a CBF for the tightened state constraint $x\in X_{\lambda}$ and $\Delta\leq\lambda$, for any initial state $x_{0}\in\mathcal{S}_{B_{k}}$ that makes problem (15) recursively feasible, the closed-loop system $x_{t+1}=f\left(x_{t},\pi_{\theta}(x_{t})\right)$ always satisfies the state constraint $x\in X$, i.e., $\max_{t\in\mathbb{I}(\infty)}\{h(x_{t})\}\leq 0$. (iii) If $B_{k}$ is a $\lambda$-contractive CBF for the original state constraint and $\Delta\leq\lambda/2$, problem (15) will be recursively feasible for all initial states in $\mathcal{S}_{B_{k}}$ and the closed-loop system $x_{t+1}=f\left(x_{t},\pi_{\theta}(x_{t})\right)$ always satisfies the state constraint $x\in X$, i.e., $\max_{t\in\mathbb{I}(\infty)}\{h(x_{t})\}\leq 0$.
In other words, in Option 2 of (V-A), we get a CBF $B_{k}$ for the tightened constraint $x\in X_{\lambda}$. The policy $\pi_{\theta}$ will make the system always satisfy the original state constraint if problem (15) is always feasible. However, the feasibility of problem (15) is only guaranteed for the initial state, i.e., problem (15) can be infeasible for some subsequent states. Practically, ones need to perform offline some statistical or deterministic verification [25] to analyze the recursive feasibility of (15). Otherwise, as recursive feasibility of (15) cannot be predicted in advance, one needs to do some relaxation for the constraint in (15) once infeasibility is observed. In comparison, in Option 3 of (V-A), we get a $\lambda$-contractive CBF $B_{k}$. An induced feature is the recursive feasibility of problem (15). As a result, under the sufficiently small approximation error, the state-action CBF approximation $Q_{\theta}$ will render $\pi_{\theta}$ a safe policy according to Definition 2.
V-D Discussion on state-action CBFs
One limitation of the parameterization (7) is that it sacrifices the universal approximation property [28] of NNs, since it restricts $Q_{\theta}$ to be quadratic on $u$. This may result in a large error bound $\Delta$ and cause constraint violations according to Theorem 2. As our main motivation for introducing state-action CBFs is to improve online computational efficiency, such a sacrifice is acceptable. Conversely, fully parameterizing $Q_{\theta}$ with a continuous NN, possessing universal approximation capabilities, increases the online computational complexity and introduces challenges in finding the global optimum of (15). In conclusion, when applying an approximation to the state-action CBF, there is typically a natural trade-off between online computational efficiency and safety.
Our proposed concept of state-action CBFs is inspired by the definition of state-action (Q) value functions in reinforcement learning [29]. Both kinds of functions do not only characterize the energy of each state but also quantify the quality of taking each action in each state. This similarity indicates that it is possible to use reinforcement learning methods [30] such as Q value iteration and Q learning to learn state-action CBFs directly, circumventing the collection of CBF samples. This methodology is plausible as the optimal control problem (V-A) can be approached via dynamic programming [30]. However, we do not adopt RL methods to learn state-action CBFs because the approximation error could accumulate when the iteration progress goes on. To interpret this, as $\alpha\geq 1$ in (V-A), $\alpha$ leads to a non-contractive dynamic programming model and consequently, convergence of RL algorithms is in general not guaranteed.
VI Case study
TABLE I: Performance of different controllers.
Safety filters Basic control policies
Learning MPC ADP LQR
Safety rate CPU time Safety rate CPU time Safety rate CPU time
No filter 63.52% 0.024 ms 36.66% 0.024 ms 56.99% 0.023 ms
Standard CBF 63.52% 2.3 ms 39.93% 2.4 ms 57.17% 2.3 ms
Quadratic state-action CBF 70.05% 1 ms 47.19% 1.1 ms 65.52% 0.9 ms
Nonlinear state-action CBF 65.15% 2.5 ms 40.19% 3.0 ms 60.80% 2.9 ms
Quadratic tightened state-action CBF 98.73% 1.0 ms 98.73% 1.0 ms 92.92% 1.0 ms
Nonlinear tightened state-action CBF 82.40% 2.7 ms 82.03% 2.7 ms 90.56% 3.0 ms
Quadratic contractive state-action CBF 100% 0.9 ms 100% 1.0 ms 100% 0.9 ms
Nonlinear contractive state-action CBF 96.91% 2.1 ms 92.20% 2.0 ms 94.57% 3.0 ms
MPC Horizon=5 Horizon=6 Horizon=7
70.24% 38.2 ms 95.83% 37.6 ms 100% 38.5 ms
We validate the proposed methods for an active inverted pendulum interacting with elastic walls, which is a piecewise affine system [3]. The detailed model can be found in Appendix -E. The simulations are conducted in MATLAB R2021a on an AMD Core R7-5800H CPU @3.20GHz machine. All optimization problems involved in the policy filters (II-B) and (15) are solved using the Matlab function “fmincon”. The problem (V-A) and the MPC problem are transformed equivalently into mixed-integer quadratic programming problems and are solved by Gurobi [31].
The system contains the state $x=[\theta\;\dot{\theta}]^{T}$ where $\theta$ is the pendulum angle, and the input $u$, which is the torque acting on the bottom end of the pendulum. The state and input constraints are given by $|x_{1}|\leq 0.15$, $|x_{2}|\leq 1$, and $|u|\leq 4$. The system is discretized using the explicit Euler method with time step 0.05s.
Using the approach in Appendix -D and letting $\lambda=0.2$ for Option 2 and $\lambda=0.05$ for Option 3, we obtain the initial CBF $B_{0}(x)=x^{T}Px-1$ with $P=\left[\begin{array}[]{rr}124.74&7.44\\ 7.44&2.24\end{array}\right]$, $\left[\begin{array}[]{rr}283.40&20.25\\ 20.25&4.69\end{array}\right]$, and $\left[\begin{array}[]{rr}522.39&25.69\\ 25.69&5.39\end{array}\right]$ for Options 1, 2, and 3, respectively. We sample from the state-input space $\left\{(x,u)\big{|}\;|x_{1}|\leq 0.16,|x_{2}|\leq 1.1,|u|\leq 4\right\}$ on a uniform grid and we obtain $40^{3}$ samples $\{(x_{i}^{\mathrm{s}},u_{i}^{\mathrm{s}})\}^{40^{3}}_{i=1}$. The successor state $f\left(x_{i}^{\mathrm{s}},u_{i}^{\mathrm{s}}\right)$ of each sample $\left(x_{i}^{s},u_{i}^{s}\right)$ is fed to the CBF generator (V-A) with $k=7$. According to Section V-B, only those samples satisfying $B_{k}(f(x_{i}^{\mathrm{s}},u_{i}^{\mathrm{s}}))\leq 10$ are obtained. After this procedure, 42109, 38609 and 37027 samples are collected in each option.
We compare state-action CBFs and standard CBFs. We consider 4 different kinds of safety certificates: (i) State-action CBF for the original state constraint; (ii) State-action CBF for the tightened state constraint $|x_{1}|\leq 0.14$, $|x_{2}|\leq 0.9$; (iii) Contractive state-action CBF; (iv) Standard CBF for the original state constraint. For the learning, we use NNs with three hidden hyperbolic tangent layers containing 16, 64 and 8 neurons. We parameterize each standard CBF by an NN. For the parameterization of each state-action CBF, we use (7) or fully parameterize it by an NN. Besides, we also consider 4 different control policies. The first three include (i) supervised learning the MPC policy [19], (ii) the approximate dynamic programming (ADP) policy [3], which is one kind of RL policy, and (iii) the LQR policy for the linearized system around the origin. These policies may be unsafe and are thus taken as $\pi_{0}$ in (II-B) and (15). The learning-based control policies (i) and (ii) are also represented by NNs. All NNs are trained via the stochastic gradient descent algorithm [32]. The fourth one is implicit MPC. One can refer to Appendix -E for a detailed description of these policies.
To test all the control methods, we randomly sample initial states from the state sample set $\{x_{i}^{\mathrm{s}}\}^{40^{3}}_{i=1}$ and select a total of 551 initial states $x$ that satisfy $B_{k}(x)\in(-0.1,0)$. These initial states are contained in the safe set of $B_{k}$, and thus can be steered to the origin without constraint violation. Meanwhile, since the values of $B_{k}$ taken at these states are close to zero, these initial states are near the boundary of the safe set. Therefore, they are in general more difficult to stabilize than any other states in the safe set. For each initial state, we simulate the closed-loop systems for 50 time steps.
The simulation results are shown in Table 1. The performance metrics include the rate of safety and the average CPU time for computing the control input per time step. The safety rate is defined as the number of initial states leading to a safe closed-loop trajectory divided by the total number (551) of initial states. Table 1 shows that directly applying the basic policies yields a lower rate of constraint satisfaction compared to passing these policies through safety filters. Among all types of CBFs, contractive state-action CBFs perform the best in the rate of constraint satisfaction, resulting in almost zero constraint violations. The superior performance of the quadratic contractive state-action CBF over its nonlinear counterpart with respect to constraint satisfaction is attributed to the emergence of suboptimal solutions to (15) when employing the nonlinear contractive state-action CBF. State-action CBFs with state constraint tightening make a significant improvement in safety performance compared to those CBFs without constraint tightening. Regarding the CPU time, we find that the implicit MPC controllers need the most time to compute the control inputs. Standard CBFs and state-action CBFs with a full NN parameterization have a comparable online computation time (from 2 to 3 ms per time step), while state-action CBFs with the parameterization (7) achieve a dramatic decrease to near 1 ms per time step. It should be noticed that the computational advantage of the quadratic state-action CBF will be larger if multi-start optimization is used to solve (II-B) or (15) with a fully parameterized CBF.
VII Conclusions and future work
This paper has presented a new type of CBFs, called state-action CBFs, which can synthesize safe control policies for general constrained nonlinear systems with a very small online computational overhead. The paper developed a new approach to learn the proposed CBF from data. We also proposed a constraint tightening approach to improve the robustness of the learned CBFs to learning errors. Simulation on an inverted pendulum with hybrid dynamics shows that the proposed method can achieve zero constraint violations and can compute the control input in about 1 ms.
In the future, we will first focus on learning the proposed CBFs for large-scale problems in a computationally efficient manner. We will also investigate how the proposed CBFs can be exploited to guide the learning of the basic policy. This will help to reduce the potential degradation of other control performance measure caused by the safety filter. More importantly, as the proposed safety filter does not need any information about the model, it can be applied in model-free cases provided that the state-action CBFs can be learned without knowing the model. As model-free (reinforcement) learning control has received much attention recently, the state-action CBF approach is an important future research direction to achieve safe learning-based control in a completely model-free manner.